five years from now, the sum of the ages of a woman and her daughter will be 40 years. the difference in their present age is 24 years. how old is the daughter
Answer by stanbon(75887)

five years from now, the sum of the ages of a woman and her daughter will be 40 years. the difference in their present age is 24 years. how old is the daughter
---------------------------------
Equaton::
m+5 + d+5 = 40
m - d = 24
--------
Rearrange:
m + d = 30
m - d = 24
-----
Add and solve for "m":
2m = 54
Mother's age = 27 yrs
-----
Solve for "d":
m + d = 30
27 +d = 30
Daughter's age = 3 yrs.
==========================
Cheers,
Stan H.
====================